# SparseLLM: Towards Global Pruning of Pre-trained Language Models

### v3

@@ A.5 Experiment Results for Additional Models
OPT-13b
Sparsity Method BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
Dense 65.87 57.76 52.44 66.02 67.82 33.46 28.62 53.14
70% SparseGPT 63.03 54.87 50.89 65.43 67.47 32.85 26.40 51.56
SparseLLM 63.85 55.23 50.73 65.67 66.46 31.83 27.20 51.57
80% SparseGPT 59.72 52.35 46.82 61.48 62.50 31.23 21.80 47.99
SparseLLM 60.89 53.07 46.19 62.12 62.21 30.38 23.00 48.27
90% SparseGPT 47.49 52.71 33.17 51.54 39.98 21.33 17.80 37.72
SparseLLM 53.43 52.71 38.19 52.96 46.68 25.26 17.40 40.95
3:4 SparseGPT 47.55 53.43 31.30 50.20 37.63 22.53 17.60 37.18
SparseLLM 51.13 52.35 38.51 55.96 49.24 24.83 21.40 41.92

### v4

@@ A.5 Experiment Results for Additional Models
A.6 Hyperparameter Selection
Hyperparameters $\alpha$ and $\beta$ are used in Eq. 5. We select $\alpha$ and $\beta$ from the set $\{0.01,0.1,1,5,10,100\}$ and perform a study on models to understand the impact of the hyperparameters. Results for OPT-1.3b with 70% sparsity are shown in Table 10.
Table 10: Ablations of the hyperparameters $\alpha$ and $\beta$ on OPT-1.3b with 70% sparsity (in perplexity)
$\alpha$ / $\beta$ $0.01$ $0.1$ $1$ $5$ $10$ $100$
$0.01$ 18.01 17.97 17.97 - - -
$0.1$ 18.04 17.82 17.96 18.04 18.40 -
$1$ 18.20 18.02 18.11 17.87 17.96 18.22
$5$ 18.06 18.02 18.03 17.92 17.96 18.04
$10$ 18.03 18.01 17.96 17.96 17.96 18.03
$100$ 18.04 18.04 17.98 18.01 18.01 18.03
A.7 Limitations and Future Work
While SparseLLM marks a significant step forward in the efficient pruning of large language models, it is important to acknowledge the inherent trade-offs associated with any model compression technique. Firstly, while our method reduces the complexity of LLMs and enhances computational efficiency, there is an inevitable balance between sparsity and performance that requires careful calibration. Additionally, in this work, we still assume homogeneous sparsity, i.e., the pruning sparsity for each layer is the same and equal to the global sparsity. How to achieve heterogeneous sparsity under our framework and fully fulfill the potential of global pruning is of great interest. Lastly, the effectiveness of SparseLLM, like any pruning method, may vary across different models and tasks, and its generalizability to all scenarios remains an area for further exploration.